OpenStudy (anonymous):

Find the Mass of both objects M+m=2.668e18 where the bigger m is 1.74 bigger. need help finding the mass of each m

OpenStudy (anonymous):

If M=1.74m then, \[1.74m+m=2.668*10^{18}\]Solve for m. Once m is known, use M=1.74m to get M.

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

first you are given that M is 1.74 times bigger than m. This means that M=1.74*m You are also given that M+m=2.668e18 Since M=1.74m, we put this in the above equation and get 1.74m+m=2.668e18 this gives us, 2.74m=2.668e18 I assume you can solve this from here.

OpenStudy (anonymous):

yes i have done that but it still comes out wrong for some reason

OpenStudy (anonymous):

63934597.69 is what i get

OpenStudy (anonymous):

I get: \[m=\frac{2.668*10^{18}}{2.74}=9.74*10^{17}\]Is this what you got?

OpenStudy (anonymous):

o thanks ive been miscalculating

OpenStudy (anonymous):

np

OpenStudy (anonymous):

this is a stupid question but how do i get the little m just subtration

OpenStudy (anonymous):

use the other formula. since M=1.74m and when know m=9.74e17, just carry out the multiplication. M=1.74*9.74e17

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

no prob